OpenStudy (anonymous):

show the product 1 1/6 x 2 1/2 on the rectangle

563blackghost (563blackghost):

\(\huge{\frac{1}{2}}\) is \(\huge{\frac{3}{6}}\)...So we can mutliply... \(\huge\color{red}{1\frac{1}{6} \times 2 \frac{3}{6}}\)